OpenStudy (anonymous):

Simplify: 4 6/7 + 9 1/4 + 2 3/4 + 1/7.(Points : 1) 19 18 17

OpenStudy (angela210793):

35/7+37/4+11/4=35/7+48/4=35/7+12=17

OpenStudy (anonymous):

17 is the good answer